# The influence of substantial intragranular orientation gradients on the micromechanical response of heavily-worked material

### v1

@@ 5 Discussion
Above, we presented results examining variations in micromechanical response and slip system activation found when polycrystals contain intragranular orientation gradients that are often present in wrought alloys after thermomechanical processing. These simulations were made possible by novel HEDM reconstruction and mesh instantiation methodologies. We found that in the Al-7085 studied, the incorporation of intragranular orientation gradients leads directly to increased localization of stress, decreased localization of plastic strain, and increases in the number of slip systems activated to accommodate deformation. We interpreted the differences in deformation response between these two microstructures as indication that neglecting intragranular orientation gradients in simulation instantiation (i.e., the typical practice) can lead to significantly different micromechanical responses from the true material being modeled.
We believe that neglect of these effects is likely not from lack of awareness, but rather that measurement of 3D orientation fields at sub-grain length scales has been historically out of reach. With maturation of 3D high-energy X-ray diffraction and 3D EBSD measurement capabilities (both of which now have commercial products available), consideration of 3D orientation fields for modeling purposes is increasingly feasible. For the remainder of the discussion, we will discuss the generality of the modeling results presented in addition to implications for modeling of the fatigue life of alloy systems.
@@ 5.1 Role of Elastic Anisotropy
Aluminum alloys are relatively elastically isotropic, with Zener ratios close to 1. The Zener ratio for the elastic moduli used in all simulations presented thus far is 0.94, which led to very similar distributions of von Mises stress between the virtual microstructures in the elastic regime (see Figure 7a), despite the difference in microtexture. To evaluate the role that elastic anisotropy would have on the micromechanical response, we perform a second simulation on the gradient microstructure, but with modified elastic properties delimited in the second row of Table 1. Here, we modified the elastic moduli to increase the Zener ratio (to 2.98, or approximately 3 that of the original moduli) while maintaining the bulk modulus, which is mutually achieved via an increase in . A Zener ratio near 3 is similar to what is found in the relatively highly elastically anisotropic systems of Fe, Cu, and Ni cubic alloys [36]. We will refer to simulations using the elastic moduli in the top and bottom rows of Table 1 as “isotropic” and “anisotropic” simulations, respectively. We note that the increase of shear modulus for the anisotropic moduli also had the effect of increasing the total stiffness of the microstructure.
@@ 5.2 Effect of Orientation Gradients on Fatigue Indicator Parameters
where in all cases denotes the slip system, is the difference in slip system shear strains at the cyclic end points, is the maximum normal stress on a slip system, is the macroscopic yield strength (see Table 3), and is a material-dependent weighting of the normal stresses. The value of generally ranges from 0.5 to 1, but was found to be as high as 10 for Al-7075. Here, we followed this procedure and cyclically loaded both the homogenized and the gradient microstructures to macroscopic strain end points of for 3 cycles, and followed the approach of Stopka et al. for the weighting of the normal stresses (i.e., ). We depict the macroscopic cyclic response of the homogenized (blue) and gradient (red dashed) microstructures in Figure 13a. We observe cyclic hardening in the response of both microstructures, with the cyclic hardening being larger in the gradient microstructure—consistent with the increased hardening rate observed during the uniaxial loading.
@@
As discussed in the previous subsection, anisotropy of elastic response has the largest effect in the elastic regime and in the elasto-plastic transition. As such, it is plausible that elastic anisotropy will play a role during cyclic loading, assuming the strain states are not appreciably large so as to diminish the influence of elastic anisotropy, as demonstrated previously. For this reason, we repeated the cyclic loading simulations and FIP analysis for the homogenized and gradient microstructures, but using the previously-introduced anisotropic elastic moduli. We present the macroscopic cyclic response for the gradient and homogenized microstructures in Figure 13b, as well as the element-by-element comparison of the maximum slip system in Figure 14b. In both microstructures, we observe a moderate increase in the amount of cyclic hardening when considering the anisotropic moduli. Additionally, with increased anisotropy (and stiffness), we note that the maximum values in both microstructures increases (in addition to the previously-observed trend of the gradient structure exhibiting higher values). From these comparisons, we interpret the trends to indicate that increased elastic anisotropy magnifies the effects of intragranular orientation gradients, and will further lower fatigue life in comparison to an elastically isotropic alloy. However, we note that there is less difference in the histograms for the anisotropic simulations compared to the isotropic simulations, i.e., both simulations display maximum FIP values of similar magnitude, indicating their maximum likelihood of fatigue to be similar. Statistics from the histograms are again presented in Table 6, which here, instead, show that the homogenized microstructure now has a higher volume fraction with extreme FIP values. In light of the similar maximum FIP values in both microstructures, we now give more weight to the difference in volume fractions to indicate fatigue performance, which thus again indicates that the homogenized microstructure would fail earlier than the gradient microstructure.
@@ 6 Summary
In this study, we utilized HEDM characterization of a 3D orientation field in forged Al-7085 to explore the role of intragranular orientation heterogeneity on both macroscopic and microscopic mechanical deformation response. We performed HEDM experiments in which we measured a 3D spatial orientation field—containing both low-angle boundaries and large amounts of intragranular misorientation ()—which we utilized to instantiate virtual microstructures with (and without) intragranular orientation heterogeneity. We utilized crystal plasticity finite element simulations to predict the deformation response of these two microstructures subjected to both monotonic and cyclic loading, and additionally altered the elastic parameters to explore the influence of elastic anisotropy on the observed trends as an effort to lend insight to expected behavior in other material systems. From the simulation results, we find that, relative to a microstructure with homogenized grain orientations, the incorporation of intragranular orientation gradients:
@@
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.

### v2

@@ 5 Discussion
Above, we present results examining variations in micromechanical response and slip system activation found when polycrystals contain intragranular orientation gradients that are often present in wrought alloys after thermomechanical processing. These simulations were made possible by novel HEDM reconstruction and mesh instantiation methodologies. We find that in the Al-7085 studied, the incorporation of intragranular orientation gradients leads directly to increased localization of stress, decreased localization of plastic strain, and increases in the number of slip systems activated to accommodate deformation. We interpret the differences in deformation response between these two microstructures as indication that neglecting intragranular orientation gradients in simulation instantiation (i.e., the typical practice) can lead to significantly different micromechanical responses from the true material being modeled. While there were only minor differences in the macroscopic stress-strain response (primarily hardening rates), the locations of extreme events (both stress and plastic deformation activity) shifted significantly between the two microstructure which in turn will lead to differences in failure locations and properties such as fatigue life.
We believe that neglect of these effects is likely not from lack of awareness, but rather that measurement of 3D orientation fields at sub-grain length scales has been historically out of reach. With maturation of 3D high-energy X-ray diffraction and 3D EBSD measurement capabilities (both of which now have commercial products available), consideration of 3D orientation fields for modeling purposes is increasingly feasible. Importantly, we believe that details of microstructural representation, such as intragranular orientation gradients, will be a key contributor to translating failure predictions from statistical to deterministic efforts. For the remainder of the discussion, we explore the generality of the modeling results presented in addition to implications for modeling of the fatigue life of alloy systems.
@@ 5.1 Role of Elastic Anisotropy
Aluminum alloys are relatively elastically isotropic, with Zener ratios close to 1. The Zener ratio for the elastic moduli used in all simulations presented thus far is 0.94, which led to very similar distributions of von Mises stress between the virtual microstructures in the elastic regime (see Figure 7a), despite the difference in microtexture. To evaluate the role that elastic anisotropy would have on the micromechanical response, we perform a second simulation on the gradient microstructure, but with modified elastic properties delimited in the second row of Table 1. Here, we modify the elastic moduli to increase the Zener ratio (to 2.98, or approximately 3 that of the original moduli) while maintaining the bulk modulus, which is mutually achieved via an increase in . A Zener ratio near 3 is representative of a vast majority of FCC alloys including stainless steels, Cu alloys, and Ni-based superalloys [36]. We refer to simulations using the elastic moduli in the top and bottom rows of Table 1 as “isotropic” and “anisotropic” simulations, respectively. We note that the increase of shear modulus for the anisotropic moduli also has the effect of increasing the total stiffness of the microstructure.
@@ 5.2 Effect of Orientation Gradients on Fatigue Indicator Parameters
where in all cases denotes the slip system, is the difference in slip system shear strains at the cyclic end points, is the maximum normal stress on a slip system, is the macroscopic yield strength (see Table 3), and is a material-dependent weighting of the normal stresses. The value of generally ranges from 0.5 to 1, but was found to be as high as 10 for Al-7075. Here, we follow this procedure and cyclically load both the homogenized and the gradient microstructures to macroscopic strain end points of for 3 cycles, and follow the approach of Stopka et al. for the weighting of the normal stresses (i.e., ). We depict the macroscopic cyclic response of the homogenized (blue) and gradient (red dashed) microstructures in Figure 13a. We observe cyclic hardening in the response of both microstructures, with the cyclic hardening being larger in the gradient microstructure—consistent with the increased hardening rate observed during the uniaxial loading.
@@
As discussed in the previous subsection, anisotropy of elastic response has the largest effect in the elastic regime and in the elasto-plastic transition. As such, it is plausible that elastic anisotropy will play a role during cyclic loading, assuming the strain states are not appreciably large so as to diminish the influence of elastic anisotropy, as demonstrated previously. For this reason, we repeat the cyclic loading simulations and FIP analysis for the homogenized and gradient microstructures, but using the previously-introduced anisotropic elastic moduli. We present the macroscopic cyclic response for the gradient and homogenized microstructures in Figure 13b, as well as the element-by-element comparison of the maximum slip system in Figure 14b. In both microstructures, we observe a moderate increase in the amount of cyclic hardening when considering the anisotropic moduli. Additionally, with increased anisotropy (and stiffness), we note that the maximum values in both microstructures increases (in addition to the previously-observed trend of the gradient structure exhibiting higher values). From these comparisons, we interpret the trends to indicate that increased elastic anisotropy magnifies the effects of intragranular orientation gradients, and will further lower fatigue life in comparison to an elastically isotropic alloy. However, we note that there is less difference in the histograms for the anisotropic simulations compared to the isotropic simulations, i.e., both simulations display maximum FIP values of similar magnitude, indicating their maximum likelihood of fatigue to be similar. Statistics from the histograms are again presented in Table 6, which here, instead, show that the homogenized microstructure now has a higher volume fraction with extreme FIP values. In light of the similar maximum FIP values in both microstructures, we now give more weight to the difference in volume fractions to indicate fatigue performance, which thus again indicates that the homogenized microstructure would fail earlier than the gradient microstructure.
@@ 6 Summary
In this study, we have presented a novel method to reconstruct 3D orientation fields with large amounts of intragranular misorientation. We demonstrate the approach through reconstruction of a forged Al-7085 specimen—containing both low-angle boundaries and large amounts of intragranular misorientation ()— and utilize the reconstructed sample to explore the role of intragranular orientation heterogeneity on both macroscopic and microscopic mechanical deformation response. We utilized crystal plasticity finite element simulations to predict the deformation response of these two microstructures subjected to both monotonic and cyclic loading, and additionally altered the elastic parameters to explore the influence of elastic anisotropy on the observed trends as an effort to lend insight to expected behavior in other material systems. From the simulation results, we found that—relative to a microstructure with homogenized grain orientations—the incorporation of intragranular orientation gradients:
@@
Data Availability Statement
The raw/processed data required to reproduce these findings cannot be shared publicly at this time due to technical or time limitations. Data can be made available upon reasonable request.